OpenStudy (anonymous):

Medals medals medals! the polar coordinates of a point are given. Find the rectangular coordinates of each point... 1. (2, 3pi/4) 2. (-4, 7pi/6) 3. (2/3, -2pi/3)

OpenStudy (anonymous):

@dan815 @Compassionate

hartnn (hartnn):

\(\Large x= r\cos \theta \) \(\Large y= r\sin \theta \)

hartnn (hartnn):

so if you're trying first one x= 2 cos (3pi/4) =... y= 2 sin (3pi/4) = ..

OpenStudy (anonymous):

okay, let me work this out (thanks!!) one sec

OpenStudy (anonymous):

\[-\sqrt2, \sqrt2\] ?

hartnn (hartnn):

\(\huge \checkmark \)

hartnn (hartnn):

same formula for others

OpenStudy (anonymous):

awesome! Thank you! okay okay correct me if im wrong, the next one is \[-\sqrt3/2, 2\]

hartnn (hartnn):

2 is correct... are you sure -sqrt 3/2 is correct??

OpenStudy (anonymous):

no, i'll trying again but im stuck. Im not exactly the go to person on math so i move pretty slow

hartnn (hartnn):

cos 7pi/6 is indeed -sqrt 3/2 but x is actually r cos theta = -4 * (-sqrt 3/2) = 2sqrt 3 :)

OpenStudy (anonymous):

alright, thats clear! thanks

hartnn (hartnn):

tell me what u get for the last one...

OpenStudy (anonymous):

alright, brace yourself

hartnn (hartnn):

*fingers crossed*

OpenStudy (anonymous):

\[(-4/3, 4\sqrt3/3)\]

hartnn (hartnn):

what did u take as 'r' ?

OpenStudy (anonymous):

2/3?

hartnn (hartnn):

and cos (-2pi/3) = -1/2 so x= (2/3)*(-1/2) = ... ?

hartnn (hartnn):

i guess you divided by -1/2 instead of multiplying :3

OpenStudy (anonymous):

comman misconception >.>

hartnn (hartnn):

whats your new co-ordinates ?

OpenStudy (anonymous):

1/3 ?

hartnn (hartnn):

x= (2/3)*(-1/2) = -1/3 right ? what about y ?

OpenStudy (anonymous):

\[\sqrt3/3\]

hartnn (hartnn):

you got the negative sign again :P \(y=-\sqrt3/3\)

OpenStudy (anonymous):

alright, that makes a lot more sense. So i can actually see where my math went wrong. Thanks a lot for clearing that up!

hartnn (hartnn):

welcome ^_^